Solve the following proportion: s/s-5=s+5/24
Answer by algebrahouse.com(1659)

Solve the following proportion: s/s-5=s+5/24

s/(s - 5) = (s + 5)/24
(s - 5)(s + 5) = 24(s) {cross-multiplied}
s^2 - 25 = 24s {multiplied through}
s^2 - 24s - 25 = 0 {subtracted 24s from both sides}
(s - 25)(s + 1) = 0 {factored into two binomials}
s - 25 = 0 or s + 1 = 0 {set each factor equal to 0}
s = 25 or s = -1 {solved each equation for s}
www.algebrahouse.com